OpenStudy (anonymous):

Given: The measure of each interior angle of a regular n-gon is x times that of an exterior angle. a.Express x in terms of n b.For what values of n will x be an integer

OpenStudy (anonymous):

The sum on an int. angle and ext. angle is 180 deg; The sum of all int. angles in a regular n-gon is 180 * (n-2); ( = 180n - 360) Therefore one int. angle in a n-gon is (180n - 360)/ n. (because a regular n-gon has n angles) ( = 180 - 360/n) One ext. angle of the same n-gon is 180 - (180 - 360/ n) ( = 360/n) x = (180 -360/n)/(360/n) x = n/2 -1 for x to be an integer, n/2 -1 must be an integer; Thus following the logic n/2 is an integer too. Therefore n is any integer times 2; n is even.

OpenStudy (anonymous):

Forgot to add that n must be bigger than 2 as 2-sided polygon does not exist in euclidean geometry.

OpenStudy (anonymous):

Thank you so much! For some reason I got (n-2)/2=x

OpenStudy (anonymous):

I did (180n-360)n = x (360/n) and then I must of messed things up